OpenStudy (anonymous):

Solve the equation, check for extraneous solutions 3x^(3/4)=192

OpenStudy (anonymous):

divide by 3 to get \[\sqrt[4]{x^3}=64\] take the cube root of both sides to get \[\sqrt[4]{x}=4\] and so \[x=4^4=256\]